OpenStudy (anonymous):

need help factoring x^2+6xy+9y^2-1

OpenStudy (anonymous):

x2+6xy+9y2−1 Since both terms are perfect 2th roots, find the values a=(x+3y) and b=1. (((x+3y))2−(1)2) The binomial can be factored using the difference of squares formula, because both terms are perfect squares. The difference of squares formula is a2−b2=(a−b)(a+b) where a=(x+3y) and b=1. ((x+3y)−1)((x+3y)+1) Simplify inside each of the factors.